HOW MUCH SALT MUST BE ADDED TO 60 KG OF A 20% SOLUTION OF SALT TO INCREASE IT TO 40% SOLUTION
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of salt added. Note: that is a 100% solution of salt.
.20*60kg + 1.00x = .40(60kg + x)***
solve for x
.6x = .20*60kg
x = 20 kg of salt needs to be added
checking out answer
12 + 20 = .40*80 = 32